a class has 1 more boy than girl one third of the boys and three fourths of the girls love pizza if 9 students love pizza how many girls and how many boys are in the class
Answer by mananth(16949)

let number of girls be x
boys are x+1
..
(x+1)/3 + 3x/4 =9
LCD = 12
(4*(x+1) +3*3x) /12 = 9
4x+4+9x=12*9
13x+4=108
-4
13x+4-4=108-4
13x=104
x= 8 number of girls
x+1 = boys = 8+1 =9